# u-µP: The Unit-Scaled Maximal Update Parametrization

### v1

@@ Abstract
The Maximal Update Parametrization (µP) aims to make the optimal hyperparameters (HPs) of a model independent of its size, allowing them to be swept using a cheap proxy model rather than the full-size target model. We present a new scheme, u-µP, which improves upon µP by combining it with Unit Scaling, a method for designing models that makes them easy to train in low-precision. The two techniques have a natural affinity: µP ensures that the scale of activations is independent of model size, and Unit Scaling ensures that activations, weights and gradients begin training with a scale of one. This synthesis opens the door to a simpler scheme, whose default values are near-optimal. This in turn facilitates a more efficient sweeping strategy, with u-µP models reaching a lower loss than comparable µP models and working out-of-the-box in FP8.
@@ 1.1 Contributions
    Out-of-the-box FP8 training: u-µP models generate tensors that lie close to the center of a floating point format’s range, meaning that most matrix multiplications can be performed in FP8 via a simple .to(float8) cast, with a minority requiring lightweight dynamic scaling for large-scale training (Section 4.2).
@@ 3.1 Not all training setups give µTransfer
We answer this Figure 2. The first training setup (a) is aligned with that used in Tensor Programs V (their Figure 4). There are several atypical aspects to their training setup, primarily the use of a constant LR schedule and a high number of epochs. This overfitting regime makes validation loss unusable, and transfer misleadingly good. When we remove these and shift to a standard Llama training setup (b), optimal HPs begin to drift with width. This confirms Lingle’s findings that standard µP is in fact a poor fit for modern LM training. We fix this (c) by the removal of parameters from LayerNorms/RMSNorms, as suggested by Lingle, and the introduction of independent weight decay for AdamW, as suggested by Wortsman et al. [18]  [^2] Lingle suggests independent weight decay is unstable, but we find it to be more so than Adam or standard AdamW.. With these changes adopted, we recover the strong transfer shown in Tensor Programs V’s experiments.
@@ 4.1 Combining µP with Unit Scaling
The final change we make is to the input scaling rule, which we show in Section 4.4 is more effective if $A_{W}\leftarrow 1$ is replaced with $A_{W}\leftarrow 1/\sqrt{\mathrm{fan{\text{-}}out}}$  [^3] This represents a slight deviation from the Maximal Update Parametrization, though we still refer to our scheme as a form of µP as it conforms in all other aspects.. With these changes made, we arrive at our final u-µP scheme, given in Table 2. It’s important to note that the scaling rules in this table must be combined with the standard Unit Scaling rules for other non-matmul operations. These are covered in Appendix B, and implemented in our library (see Appendix D).
@@ 4.2 Out-of-the-box low-precision training
By applying the principles of Unit Scaling to µP, u-µP gains a key feature: the easy use of low-precision number formats during training. We can attribute the difficulties µP has with low precision to the fact that it ignores constant factors (along with weight and gradient-scaling), only ensuring that activations are of order $\Theta(1)$. The stricter condition of unit scale across all tensors at initialization provides a way of leveraging µP’s rules in order to make low-precision training work. During training, most scales in the model stabilize while certain tensors exhibit scale growth (see Section 5.5) and require a separate treatment in order to be cast to lower precision. The two solutions we propose here are either to use the E5M2 format to represent the larger scales or to apply a dynamic rescaling of the matmul input (Section 5.6 and Section 5.7 respectively).
@@ 5.4 Hyperparameter transfer
5.5 Numerical properties
Figure 6 shows the RMS over all linear modules in the model, Figure 13 shows RMS on a per-tensor basis over training steps, and Figure 14 shows the effect on RMS of LR, width, depth, steps and batch size. RMS captures the larger of the mean and scale of a distribution, and as such can be a good test of whether a tensor is likely to suffer over/underflow in low-precision number formats. Detailed analysis of these statistics is given in Section A.6, with our results supporting the central thesis of Unit Scaling: that tensors are well-scaled at initialization and largely remain so across training.
Based on these conclusions we propose our primary FP8 scheme: for every matrix multiplication, we cast the input, weight and grad-output tensors to E4M3, with the exception of the inputs to FFN and self-attention final projections, which are cast to E5M2 to accommodate their growing scale. This simply requires FP8 casts to be inserted into the model, avoiding the more complex scaling techniques used for FP8 in the literature (see Appendix J). This is possible due to the numerical stability we see in our analysis of u-µP. Additional details of our primary FP8 scheme are given in Section A.7.
As we demonstrate in Section 5.7, the primary scheme needs to be slightly adapted when model size, sequence length and number of training steps increase substantially.
5.6 FP8 training at smaller scale
We now show that u-µP can indeed be trained in our primary FP8 scheme in a smaller-scale setting (in terms of training tokens, rather than model-size). We note that our aim here is a proof-of-concept that this form of low-precision training can be done without degradation, not a demonstration of improved throughput which we leave to future work. To investigate the question of the potential benefits of FP8 training, Appendix K shows results for micro-benchmarking low-precision matmuls. We find that the addition of scaling factors adds no overhead, making our u-µP modifications essentially ‘free’.
Figure 1 (c) demonstrates the application of our FP8 scheme to model-training at width 4096. We use exactly the HPs suggested by the sweep in Figure 1 (a), but transferred to the larger model-width. µP fails entirely under our FP8 scheme due to gradient underflow, reflecting the requirement for different, and likely more complex scaling scheme. In contrast, u-µP trains in FP8 with only a small increase in validation loss versus the full-precision baseline.
@@ 5.6 FP8 training at smaller scale
5.7 FP8 training at large-scale
The largest experiments in our above plots contain 1B parameters, but are trained at a smaller sequence length and number of steps than are typically used. Here we demonstrate that u-µP enables stable FP8 training for more realistic large-scale training scenarios [^4] The models in this section were trained on a separate code base optimized for distributed training that will be made available in the near future., although some tweaks to our primary FP8 scheme are required, as outlined below. We use the same architecture as before, again following the training setup recommendations in Appendix C, and train our target models on 300B tokens of the publicly available SlimPajama dataset [22] (see Section A.8 for training details). We use three target model-sizes at approximately 1B, 3B and 7B parameters.
@@ 5.7 FP8 training at large-scale
    When using a relatively small proxy model with 8 layers and a width of 512 (4 attention heads), the HP-loss landscape is rather noisy. By doubling the width we are able to discern the optimal values of the HPs more clearly.
@@
    Learning rate $\eta$ and residual attention ratio $\alpha_{\mathrm{res\text{-}attn\text{-}ratio}}$ are the most important HPs. All other HPs can be left at their default value of $1$.
@@
    The optimal values of these HPs are $\eta=2^{3.5}$ and $\alpha_{\mathrm{res\text{-}attn\text{-}ratio}}=2^{-2.0}$ and thus differ non-trivially from the observed HPs in our previous experiments.
We also note that the previously-observed scale growth of inputs to FFN and self-attention final projections increases here. Directly casting inputs to FP8 fails in this setting, including when using E5M2 for extended numerical range. All other layers can still be cast to FP8 without problems (validated up to 3B scale), though we now require gradients to be in E5M2. To tackle the layers that exhibit scale growth, we fit them with a lightweight form of dynamic rescaling, leaving the rest of the model untouched.
The dynamic rescaling works as follows: Before the matrix multiplication, we normalize the input by its standard deviation and divide the result by the same scale afterwards, while ignoring both of these operations in the backward computation. We emphasize that this is still considerably less complex than the per-tensor scaling strategy that is usually required for FP8 training.
Using this refined FP8 scheme we perform two series of experiments:
  1. 1.
    On the 1B scale, we train a comparable SP model adopting the learning rate and initialization scheme from the Pythia model family [23], which we consider a strong established baseline. Apart from a standard training in BF16 mixed precision, we also try a simple FP8 cast and the Nvidia Transformer Engine framework [17] for this baseline. We then compare these SP models to three variants of our u-µP 1B model: One model in mixed precision, one model casting all layers except the critical ones to FP8 (our ‘partial’ scheme in Figure 7), and one where we apply our full FP8 scheme. Overall the u-µP models perform favorably (see Figure 7 (left)).
  2. 2.
    We train FP8 models up to 7B parameters using our full FP8 scheme. All runs converge (see Figure 7 right), and we expect our method to work at larger scales still.
Figure 7: (Left) Comparison of 1B training curves for u-µP and SP, using BF16 mixed precision vs FP8. Directly casting to FP8 causes SP to diverge, whereas u-µP still converges and outperforms the Transformer Engine FP8 baseline (SP FP8 T.E.). Using our proposed partial FP8 scheme, u-µP maintains nearly full performance compared to BF16 mixed precision. (Right) Loss curves from large scale training runs up to 7B using the u-µP FP8 scheme. Both figures use a smoothed running average loss with window size 100.
@@ 7 Conclusions
We present an improved version of µP, underpinned by the principle of Unit Scaling. This provides the simple low-precision training that comes with unit-scale, but also provides a platform for solving other problems with µP. For instance, many HPs lack a clear interpretation under µP, but with u-µP we associate them with non-linear functions that have unit-scaled inputs, which means they can be understood as (inverse) temperatures. We build other improvements to the HP scheme on top of this, dropping the unnecessary and unhelpful base-shape HPs and $\sigma_{\mathrm{init}}$, and and minimizing dependencies between the remaining HPs.
This results in a simpler scheme to implement than µP, with several beneficial properties. A more efficient independent search can be done for proxy model HPs under u-µP, which can even drop non-LR HPs and still reach near-optimal loss. Our improved embedding LR scaling facilitates better performance at large widths, and we see strong HP transfer across width, depth, batch size and steps. Overall, u-µP simplifies and strengthens the practical application of µP, and provides further evidence that the principle of Unit Scaling is beneficial for model design.
@@ A.5 Hyperparameter transfer experiments
Whilst Figure 11 demonstrates the transfer of individual hyperparameters over width, Figure 12 instead demonstrates the simultaneous transfer of all hyperparameters when co-optimized on the small-scale proxy model, as is done for µTransfer. The µP and u-µP points are taken from Figure 1, with hyperparameters swept on a model of width 256 using a full random HP search and a simple learning rate sweep for µP and u-µP respectively. The Standard Parametrization scheme, as shown in Table 3 requires choosing a learning rate and a weight-initialization scheme. We follow the initialization scheme of GPT-NeoX [48], and transfer learning rate using a heuristic scaling factor of $\nicefrac{{\mathrm{base{\text{-}}width}}}{{\mathrm{width}}}$.
Figure 12: Transferring hyperparameters from width 256 up to 4096 using three different hyperparametrization schemes. µP and u-µP results are as seen in Figure 1, whilst Standard Parametrization follows the initialization approach of GPT-NeoX [48].
@@ A.6 Numerical properties
Figure 14: The effect of hyperparameters on FP8 training loss and on the end-training RMS of various tensors: (a) decoder weight, (b) last-layer FFN down-projection input and (c) last-layer FFN down-projection output gradient. Only learning rate has a substantial effect on the end-training RMS. Vertical lines show the default setting of that hyperparameter, as used for all other plots.
A.7 Our primary FP8 scheme
Each linear module in a model induces three matrix-matrix products: one during the forward pass to compute the output and two during the backward pass, computing gradients w.r.t. the weights and the inputs, respectively. The tensors participating in these matrix-matrix products are the input activations, the weights and the gradients w.r.t. the output activations.
We cast the input, weight and grad-output tensors to the narrower but more accurate E4M3 format, and for the two tensors which grow considerably across training—the inputs to FFN and self-attention final projections—we cast to the wider E5M2 format. The output of each matrix multiplication is produced directly in the higher-precision format (FP32 in our case), with no loss scaling or per-tensor scaling applied.
Unit scaling ensures RMS of 1 for all three tensors at initialization, and empirically the scales of these tensors do not drift too much during training, with the exception of the input tensors to the FFN and self-attention final projections, for which we use a wider format due to their increase in RMS over training steps (see Figure 13). Note that this scale growth is consistent across different HP settings (see Figure 14).
We have to adapt this scheme slightly for our large-scale training setup, which experiences increased scale-growth for the inputs to FFN and self-attention final projections, exceeding the range of E5M2. Our simple solution to this is a lightweight dynamic re-scaling for those particular problematic operations, still in FP8. This scheme is described in more detail in Section 5.7.
A.8 u-µP and FP8 at large scale
@@ A.7 Our primary FP8 scheme
We use mixed-precision during training with optimizer states in FP32 that are sharded via ZeRO stage 1 [49]. We retain the model weights in BF16. Our FP8 scheme is as follows:
  • In all matrix multiplications in the transformer block, we use E4M3 for the activation and the weight and E5M2 for the output gradient.
  • For the attention dense projection and the FFN out projection we use dynamic rescaling as described in Section 5.7.
  • In all remaining matrix multiplications we perform a simple FP8 cast with the aforementioned data types.
All other tensors remain either in BF16 (embedding, readout layer, norm, activation function) or FP32 (Flash Attention [50]).
@@ Appendix I From µP to u-µP
The default format used for training is the single-precision floating-point format, commonly known as FP32, with some hardware providers automatically casting it to the smaller TF32 compute mode for accelerated arithmetic. The 16-bit FP16 and BF16 formats were later introduced, and more recently the FP8 E5 & E4 formats [65, 66, 67]. The higher range of E5 has typically been used for gradients, while the higher precision of E4 has been seen as necessary for weights and activations. Our particular implementation of FP8 training is covered in Section A.7. Other aspects of training such as the optimizer state and cross-device communication have also been put into FP8 [27], though not all tensors are amenable to being run in the lowest precision [38] without degradation. The use of multiple formats is known as mixed precision [68]. A comparison of these formats is given in Table 10.

### v2

@@ Abstract
The Maximal Update Parametrization (µP) aims to make the optimal hyperparameters (HPs) of a model independent of its size, allowing them to be swept using a cheap proxy model rather than the full-size target model. We present a new scheme, u-µP, which improves upon µP by combining it with Unit Scaling, a method for designing models that makes them easy to train in low-precision. The two techniques have a natural affinity: µP ensures that the scale of activations is independent of model size, and Unit Scaling ensures that activations, weights and gradients begin training with a scale of one. This synthesis opens the door to a simpler scheme, whose default values are near-optimal. This in turn facilitates a more efficient sweeping strategy, with u-µP models reaching a loss that is equal to or lower than comparable µP models and working out-of-the-box in FP8.
@@ 1.1 Contributions
    Out-of-the-box FP8 training: u-µP models generate tensors that lie close to the center of a floating point format’s range, meaning that most matrix multiplications can be performed in FP8 via a simple .to(float8) cast without dynamic rescaling.
@@ 3.1 Not all training setups give µTransfer
We answer this Figure 2. The first training setup (a) is aligned with that used in Tensor Programs V (their Figure 4). There are several atypical aspects to their training setup, primarily the use of a constant LR schedule and a high number of epochs. This overfitting regime makes validation loss unusable, and transfer misleadingly good. When we remove these and shift to a standard Llama training setup (b), optimal HPs begin to drift with width. This confirms Lingle’s findings that standard µP is in fact a poor fit for modern LM training. We fix this (c) by the removal of parameters from LayerNorms/RMSNorms, as suggested by Lingle, and the introduction of independent weight decay for AdamW, as suggested by Wortsman et al. [18]  [^2] Lingle suggests independent weight decay is unstable, but we find it to be more so than Adam or standard AdamW. (see [19] for further analysis). With these changes adopted, we recover the strong transfer shown in Tensor Programs V’s experiments.
@@ 4.1 Combining µP with Unit Scaling
The final change we make is to the input LR scaling rule, which we show in Section 4.4 is more effective if $c_{W}\leftarrow 1$ is replaced with $c_{W}\leftarrow 1/\sqrt{\mathrm{fan{\text{-}}out}}$  [^3] This represents a slight deviation from the Maximal Update Parametrization, though we still refer to our scheme as a form of µP as it conforms in all other aspects.. With these changes made, we arrive at our final u-µP scheme, given in Table 2. It’s important to note that the scaling rules in this table must be combined with the standard Unit Scaling rules for other non-matmul operations. These are covered in Appendix B, and implemented in our library (see Appendix D).
@@ 4.2 Out-of-the-box low-precision training
By applying the principles of Unit Scaling to µP, u-µP gains a key feature: the easy use of low-precision number formats during training. We can attribute the difficulties µP has with low precision to the fact that it ignores constant factors (along with weight and gradient-scaling), only ensuring that activations are of order $\Theta(1)$. The stricter condition of unit scale across all tensors at initialization provides a way of leveraging µP’s rules in order to make low-precision training work.
When training a transformer model with u-µP  most scales in the model stabilize while certain tensors exhibit scale growth that potentially pushes them out of FP8 range. We empirically identify these critical tensors to be the inputs to the attention dense projection and final FFN matmul as well as the weight of the decoder head (for details see Section A.6). The latter becomes negligible in terms of model flops as width and depth of the model increase, so we generally keep this operation in higher precision.
Following these observations, we propose the following FP8 mixed precision scheme for u-µP transformer models:
  • For non-critical matmul operations, we cast the input and weight to E4M3, and the gradient with respect to the output to E5M2. This is done in the forward computation, as well as the two backward computations (for the gradient w.r.t. the weight, respectively the input). Non-critical layers are query, key, value as well as the input layer(s) to the FFN.
  • All layers involving critical tensors, as well as embedding layer, residual addition and nonlinear functions are performed in higher precision. This also means that we directly aggregate into higher precision in each FP8 matmul. We keep optimizer states in FP32, as is usually the case in mixed precision training.
We note that in some cases one can deal with the critical tensors by casting them to E5M2 instead of E4M3, however we observed some instabilities applying this in a large scale setting, possibly due to loss of precision. In small scale scenarios we also empirically find that applying the E4M3 format instead of E5M2 for the gradients is possible, but becomes problematic in a more realistic setting where gradients require a higher dynamic range.
With our proposed mixed precision scheme, about $70$% of the matmul computations in the transformer block are performed natively in FP8 (assuming a standard architecture, e.g. Llama). If desired, a dynamic per-tensor scaling could still be applied to the critical tensors.
@@ 5.4 Hyperparameter transfer
5.5 FP8 training
In this section we justify the simple mixed-precision scheme described in Section 4.2 and demonstrate that it can be used to train u-µP models out-of-the-box.
Proof-of-concept
Figure 6 shows the RMS of all linear layer inputs for a moderately sized transformer. RMS captures the larger of the mean and scale of a distribution, and as such is a good test of whether a tensor is likely to suffer over/underflow in low-precision. We observe that u-µP tensors largely have RMS starting close to $1$ and remaining so at the end of training, supporting our scheme.
Figure 14 demonstrates the scale-growth of critical tensors which our scheme is designed to accommodate, showing RMS on a per-tensor basis over steps. Figure 15 provides further insight into this issue, showing the effect of LR, width, depth, steps and batch size on the RMS of critical tensors.
As an initial proof-of-concept we train a u-µP model using our FP8 scheme over 8k steps, using HPs from a proxy model, as shown in Figure 1 (c). We see only a small degradation versus FP32, and at this scale critical tensors can still be cast to FP8 using E5M2, while gradients can even use E4M3.
@@ 5.5 FP8 training
Larger scale
Next we consider a more realistic training scenario  [^4] The training codebase used for our larger-scale experiments can be found at the following url https://github.com/Aleph-Alpha/scaling. We have also released model checkpoints, which are available at https://huggingface.co/Aleph-Alpha.. Using the same architecture, and following the steps set out in our u-µP user-guide (Appendix C), we train our target models on 300B tokens of the SlimPajama dataset [23] (see Section A.7 for training details).
@@
    When using a relatively small proxy model (8 layers and 512 width), the HP-loss landscape is rather noisy. By doubling the width we can discern optimal HP values more clearly.
@@
    The most important HPs are $\eta$ and $\alpha_{\mathrm{res\text{-}attn\text{-}ratio}}$. All others can be left at the default of $1$.
@@
    The optimal values of these HPs are $\eta=2^{3.5}$ and $\alpha_{\mathrm{res\text{-}attn\text{-}ratio}}=2^{-2.0}$ and thus differ non-trivially from the observed HPs in our smaller-scale experiments.
We then train u-µP models of approximately 1B, 3B and 7B parameters, using our FP8 mixed-precision scheme (see Section 4.2). We also train two baselines at each size: the first is a BF16 version of our u-µP models, and the second is a set of SP models using the weight init scheme from the Pythia model family [24] and the LR scheme from Llama 3 [25], scaling inversely with width and using a LR of 3e-4 at 7B scale. The loss curves are shown in Figure 7. All FP8 runs converge and show no significant loss degradation. In comparison to SP, the u-µP models have a qualitatively different training curve with a higher loss for most of training that catches up in latter stages, hinting at a fundamentally different optimization trajectory. In terms of downstream performance, both of the u-µP 7B models are competitive with SP. In particular, the scores of the FP8 model are mostly on par with the BF16 models (see LABEL:tab:eval_results).
Figure 7: Large-scale training runs. (Left) u-µP BF16 vs u-µP FP8. (Right) u-µP BF16 vs SP BF16.
Table 4: 0-shot benchmark results at 7B scale.
Scheme Format MMLU HellaSwag OpenBook QA PIQA TriviaQA WinoGr
SP BF16 29.6 52.4 27.8 76.5 22.2 63.3
u-µP BF16 29.0 53.4 31.6 77.1 23.4 63.7
u-µP FP8 31.2 53.4 29.6 77.6 21.3 65.7
@@ 7 Conclusions
We introduce u-µP, a modified and improved version of µP that satisfies Unit Scaling. Through careful analysis guided by first principles we identify an interpretable set of HPs that has minimal interdependencies and facilitates an efficient independent sweeping strategy. We show that the stability properties of µP combined with Unit Scaling enable a simple and robust FP8 mixed precision scheme that works in a realistic large scale training scenario. u-µP provides further evidence that the principle of Unit Scaling is beneficial for model design.
Limitations and future work
Some choices like the modified embedding LR rule are only justified by empirical observations, and currently lack a theoretical explanation. Additionally, neither µP nor Unit Scaling give guarantees for network quantities to be well-behaved over the course of training. In particular we would like to understand the issue (or feature) of scale growth in the critical layers better and look into possible mitigations. We also believe that low-precision training techniques can be pushed further, with u-µP offering an ideal starting point for future optimizations.
@@ A.5 Hyperparameter transfer experiments
Additionally, in Figure 11 we show learning rate transfer over sequence length for both µP and u-µP  fixing either tokens per batch or sequences per batch. In both scenarios u-µP shows not only better absolute training performance, but also better transfer behaviour as sequence length increases. Since our default proxy sequence length is 256, using µP to transfer to sequence length 2048 would result in minimal improvements or even a degradation in validation loss, whereas the u-µP shows much greater and more consistent improvements.
Figure 11: Transfer of learning rate over sequence length for µP (left) and u-µP (right). As sequence length varies, we can fix the number of tokens per batch by inversely varying the number of sequences per batch (top). Alternatively we can fix the sequences per batch and allow the number of tokens per batch to vary with sequence length (bottom). In the latter case, larger sequence lengths mean the model sees more tokens during training, though as per Table 5 this translates to >1 epoch on WikiText-103 when sequence length goes above 256.
@@
Whilst Figure 12 demonstrates the transfer of individual hyperparameters over width, Figure 13 instead demonstrates the simultaneous transfer of all hyperparameters when co-optimized on the small-scale proxy model, as is done for µTransfer. The µP and u-µP points are taken from Figure 1, with hyperparameters swept on a model of width 256 using a full random HP search and a simple learning rate sweep for µP and u-µP respectively. The Standard Parametrization scheme, as shown in Table 3 requires choosing a learning rate and a weight-initialization scheme. We follow the initialization scheme of Pythia [24], and transfer learning rate using a heuristic scaling factor of $\nicefrac{{\mathrm{base{\text{-}}width}}}{{\mathrm{width}}}$, as is done in [25].
Figure 13: Transferring hyperparameters from width 256 up to 4096 using three different hyperparametrization schemes. µP and u-µP results are as seen in Figure 1, whilst Standard Parametrization follows the initialization approach of Pythia [24].
@@ A.6 Numerical properties
Figure 15: The effect of hyperparameters on FP8 training loss and on the end-training RMS of critical tensors: (a) decoder weight, (b) last-layer FFN down-projection input and (c) last-layer FFN down-projection output gradient. Only learning rate has a substantial effect on the end-training RMS. Vertical lines show the default setting of that hyperparameter, as used for all other plots.
A.7 Large-scale training details
@@
We use mixed-precision during training with optimizer states in FP32 that are sharded via ZeRO stage 1 [50]. We retain the model weights in BF16 and apply our FP8 scheme as described in Section 4.2 to the tensors participating in matmul operations throughout the transformer block. All other tensors remain either in BF16 (embedding, readout layer, norm, activation function) or FP32 (Flash Attention [51]).
@@ Appendix I From µP to u-µP
The default format used for training is the single-precision floating-point format, commonly known as FP32, with some hardware providers automatically casting it to the smaller TF32 compute mode for accelerated arithmetic. The 16-bit FP16 and BF16 formats were later introduced, and more recently the FP8 E5 & E4 formats [66, 67, 68]. The higher range of E5 has typically been used for gradients, while the higher precision of E4 has been seen as necessary for weights and activations. Our particular implementation of FP8 training is covered in LABEL:app:fp8_scheme. Other aspects of training such as the optimizer state and cross-device communication have also been put into FP8 [29], though not all tensors are amenable to being run in the lowest precision [40] without degradation. The use of multiple formats is known as mixed precision [69]. A comparison of these formats is given in Table 11.
@@
Appendix L Attention output RMS grows with model depth
Figure 20: Scale of intermediate tensors grows with depth at initialization. Top left: Intermediate activation tensor RMS along the residual branch. Only the attention outputs after the first layer are not unit-scaled. Bottom left: Skip activation tensor RMS. Scale growth in attention outputs drives growth in skip activation scales. Note that layer_idx$=0$ corresponds to the embedding output, and layer_idx$=4$ corresponds to the final layer outputs. Top right: Intermediate gradient tensor RMS along the residual branch. Growth in the attention output scale drives growth in attention qkv gradient scales. Bottom Right: Skip gradient tensor RMS. The scale of output activations induces a global rescaling of the gradients.
A core assumption in deriving per-op scaling factors is that each input to an operation has zero mean, unit-variance, and uncorrelated elements at initialization. This is trivially true for weights and by extension the token embeddings taken as input to the transformer trunk. However, this is not guaranteed for intermediate results and gradients if an operation in the computational graph induces correlation in the elements. In such a scenario our scaling factors will not return unit-variance outputs as we will not have corrected for these correlations in the inputs. As we then increase the depth of the network, where the same operation is left to amplify correlations, we can end up with variance in intermediate results and gradients scaling with depth
Figure 20 illustrates this phenomenon in a unit-scaled four-layer Llama model with width=256. All activation tensors in the residual branches are unit-scaled, except for the output of the attention layers. We also see that the variance of attention outputs grows with depth. Since Llama models use pre-norm on the residual-branch, residual-branch inputs will revert to unit-scale again until they reach another instance of the correlation-inducing operation. As we add under-scaled attention layer results back to the skip-branch, our skip tensor variances grow with depth as our residual-add assumes unit-variance inputs. This has a knock-on effect on the global scaling of the gradients since the Jacobian of the final norm will scale the gradient by the inverse of the final skip tensor variance.
So which operation induces correlation in the attention output at initialization? For the default case where all multipliers are set to 1, our $1/d$ scaling of attention logits results in a sufficiently high temperature that attention probabilities are effectively uniform. With causal masking, we effectively take a running mean across the value tensor along the sequence dimension. As a result, each subsequent token representation is correlated with the last. Since we derive appropriate scaling factors for the first layer, we do not see scale growth emerging until the second layer, where correlations accumulate during the next effective running mean.
We leave it to future work to offer a solution to scale growth created by correlation in intermediate tensors. We note that this is scale growth emergent at initialization, but we also see scale growth in other intermediate tensors during training. Whether scale growth during training is related to the phenomenon outlined here remains to be seen.